(x-5)squared+12=0 how do I get it to the layout to pull the a b and c's out
Answer by stanbon(75887)

(x-5)squared+12=0 how do I get it to the layout to pull the a b and c's out
-----
(x-5)^2 + 12 = 0
x^2 - 10x + 25 + 12 = 0
----
x^2 -10x + 37 = 0
a = 1
b = -10
c = 37
-------
Cheers,
Stan H.
============